OpenStudy (anonymous):

Jim has a mass of 15 kilograms and is running on the ground with a velocity of 2.7 meters/second. What is Jim's kinetic energy?

OpenStudy (lgbasallote):

the formula for kinetic energy is\[\large K = \frac 12 mv^2\] mass is 15 kg and v is 2.7 m/s \[\large \implies K = \frac 12 (15)(2.7)^2\] solve that with your calculator to get your answer

OpenStudy (lgbasallote):

does that help @lycan1995 ?

OpenStudy (anonymous):

yes thank you

OpenStudy (lgbasallote):

welcome